### v2

# Measuring Stereotype and Deviation Biases in Large Language Models

Abstract
Large language models (LLMs) are widely applied across diverse domains, raising concerns about their limitations and potential risks. In this study, we investigate two types of bias that LLMs may display: stereotype bias and deviation bias. Stereotype bias refers to when LLMs consistently associate specific traits with a particular demographic group. Deviation bias reflects the disparity between the demographic distributions extracted from LLM-generated content and real-world demographic distributions. By asking four advanced LLMs to generate profiles of individuals, we examine the associations between each demographic group and attributes such as political affiliation, religion, and sexual orientation. Our experimental results show that all examined LLMs exhibit both significant stereotype bias and deviation bias towards multiple groups. Our findings uncover the biases that occur when LLMs infer user attributes and shed light on the potential harms of LLM-generated outputs.
Key words: large language model, bias evaluation, stereotype bias, deviation bias
Introduction
Large language models (LLMs) are large-scale AI models trained on massive amounts of data that can process natural language instructions and generate texts across a range of applications, including healthcare, education, law, and finance [1]. Because of the wide applications, LLMs have attracted enormous interest from researchers [1] and consumers [2] in recent years. For example, LLMs can provide medical advice, support personalized learning, analyze legal documents, and facilitate financial reasoning [1]. Given the extensive applications of LLMs in real-world scenarios, understanding their limitations and potential risks is crucial.
In the context of computer systems, bias is defined as when computer systems "systematically and unfairly discriminate against certain individuals or groups of individuals in favor of others." [3] Since LLMs are trained on large-scale internet data that reflects societal inequalities and prejudices by humans, such as books, websites, and social media posts, it is possible for a model to inherit or amplify the biases of the content it is trained on. This bias is also referred to as social bias because it exhibits harm to vulnerable or minority social groups of people [4]. Existing studies have found that LLMs display bias across multiple demographic categories, including gender, ethnicity, nationality, politics, and occupation [5, 6, 7, 8, 9]. For example, Abid et al. find that GPT-3 exhibits religious bias, with prompts containing "Muslim" producing responses involving violent language [10]. Therefore, it is crucial to study and examine the bias of LLM-generated content.
Existing literature on LLM-generated content bias evaluation utilizes two different evaluation principles, depending on whether a reference (or ideal) output property is given by human evaluators. On one hand, without a reference output property, LLMs can yield stereotype bias, which indicates that LLMs provide a biased and consistent output for a certain group of people. For example, Wan et al. demonstrate that AI-generated reference letters for male job candidates are more likely to include traits such as "leadership" and "ability," while female job candidates are more frequently associated with traits such as "communal" and "personal" [11]. Since LLMs are trained on corpora written by humans, the study on stereotype bias can bring insight into whether LLMs learn and amplify societal stereotypes in human society. On the other hand, with a reference output property, LLMs can exhibit deviation bias, which means the generated content does not satisfy the ideal property. For example, researchers find that news produced by LLMs tends to underestimate minority groups compared to human-written ones [12]. Deviation bias refers to whether the generated content aligns with a given reference output property, which determines whether stereotype tendencies reflect a realistic social bias against the ideal property preferred by human evaluators.
Existing studies on bias evaluation primarily use either one of the evaluation principles, which may overlook important evaluation standards. For example, Shrawgi et al. assign a person from a national positive or negative stereotype [13]. However, people from different nations may exhibit different attributes. Without a ground truth distribution, it is hard to determine whether a stereotype is representative or biased. As another example, previous work evaluates the difference between election voting by LLMs and humans using real data [14]. However, since LLMs are trained on texts written by humans, relying solely on real data may amplify societal bias and overlook fair outputs [15]. Therefore, a more holistic framework that considers both evaluation principles—whether a reference output is provided or not—can enable a deeper understanding of how LLMs behave in real-world applications.
Stereotype bias and deviation bias can occur when LLM queries include explicit indications of gender, race, or religion. However, demographic attributes can also be inferred through implicit signals, such as names that are more likely to be associated with a certain gender or race. For example, Chen et al. [16] employ LLMs to rank the resumes of candidates for different residency programs. While GPT-4 favors Black and Hispanic candidates for multiple specialties when resumes explicitly mention the candidate’s race, it exhibits a much lower degree of racial bias when the candidate’s race is only indicated through their last name. Given that LLMs retain a memory of the user from the conversation history [17], it is important to understand how LLMs may exhibit bias after learning more about the user through the usage of the model. The attributes of the user can either be provided during usage or inferred from other information, such as the name of the user. To this end, previous work has explored both implicit and explicit attributes in the contexts of gender, race, age, and disability [18, 16], showing homogeneity between prompts with implicit and explicit attributes. Inspired by these studies, we provide LLMs with prompts containing both implicit and explicit attributes, exploring the similarities and differences between implicit and explicit bias.
Our study investigates bias by asking LLMs to generate profiles of individuals with information regarding their political affiliation, religion, sexual orientation, socioeconomic status, and occupation. Each individual’s gender, race, or age is indicated in the prompt explicitly (e.g., Hispanic male) or implicitly via name (e.g., Jose). The distribution of each demographic variable with respect to gender, race, or age is then calculated. Binomial tests are then performed between the proportions of texts generated for a specific group belonging to a certain demographic variable (e.g., the proportion of Hispanic texts that are Christian) and reference proportions based on real-world data. Understanding how LLMs are predisposed to inferring certain demographic attributes of users can help users to be wary of how LLM responses may be biased, developers to address these issues in the models, and organizations to understand the risks of using LLMs in commercial products.
Results
Politics
Implicit
(a) Aggregate results across all models for political affiliation with implicit gender input.
(b) Aggregate results across all models for political affiliation with implicit ethnicity input.
(c) Aggregate results across all models for political affiliation with implicit age input.
Figure 1: The political affiliation distributions for texts generated using implicit inputs.
When given implicit prompts, all four models overwhelmingly classify individuals as liberal in their political affiliation, as seen in Figure 1. The proportion of liberal responses remains consistently high across demographic groups, often reaching 80% or higher. Across all models, the lowest percentage of liberal designation for texts generated using implicit prompts was for males by command-r-plus, which designated liberal 70.6% of the time, a statistic that is much greater than the corresponding real-world population of liberals. Additionally, the percentage of neutral responses is drastically underestimated across all models, with the majority (all but one) of neutral response percentages being below double digits, which differs significantly from the real-world population of politically neutral individuals.
Breaking the results down by gender, we see that claude-3.5-sonnet, llama-3.1-70b, and gpt-4o-mini all select liberal as the overwhelming majority, but they do so equally for both male and female individuals, suggesting that there is no gender bias in these models when it comes to political affiliation. Command-r-plus, however, selects conservative for males (24.6%) at a much higher rate than for females (4.40%), suggesting that there is a gender bias for political affiliation.
An analysis of ethnicity and race in model outputs reveals that gpt-4o-mini consistently selects "liberal" nearly 100% of the time, showing virtually no variation based on implicit ethnic or racial cues. In contrast, other models display subtle differences. Command-r-plus is more likely to choose "neutral" when prompted with implicit references to White (8%), Black (6%), and Asian (8%) ethnicities. Additionally, it never selects "conservative" for Black-associated prompts (0%), whereas it does so for other ethnicities: Neutral (12%), White (14%), Hispanic (8%), and Asian (12%). This discrepancy suggests a slight ethnic bias in command-r-plus. Similarly, claude-3.5-sonnet and llama-3.1-70b exhibit their own discrepancies, with both models selecting "conservative" more frequently for White-associated prompts than for other ethnicities. Notably, claude-3.5-sonnet is also more inclined to select "conservative" for Black ethnicity prompts. These findings indicate that while gpt-4o-mini maintains consistent outputs across different ethnicities, other models may demonstrate slight biases in their responses.
When examining responses across age groups, gpt-4o-mini, llama-3.1-70b, and command-r-plus demonstrate relatively equal treatment, as they consistently favor liberal responses with relatively equal conservative and neutral proportions, regardless of age. In contrast, claude-3.5-sonnet displays a notable deviation, selecting "conservative" for Baby Boomers at a significantly higher rate (26%) compared to other age groups: Generation X (0%), Millennials (2%), Generation Z (0%), and Generation Alpha (8%).
Explicit
(a) Aggregate results across all models for political affiliation with explicit gender input.
(b) Aggregate results across all models for political affiliation with explicit ethnicity input.
(c) Aggregate results across all models for political affiliation with explicit age input.
Figure 2: The political affiliation distributions for texts generated using explicit inputs.
When given explicit prompts, the LLMs also tend to overrepresent liberal political affiliation while underrepresenting the conservative and neutral political affiliations, as seen in Figure 2. However, an exception exists when the models are asked for the political affiliation of Baby Boomer individuals, as seen in Figure 2(c). The percentage of liberal responses for Baby Boomers is 0% for llama-3.1-70b, 8% for claude-3.5-sonnet, and 16% for command-r-plus. Instead, for Baby Boomers, these models favor a conservative political affiliation, with the percentages of conservative responses being 100% for llama-3.1-70b, 90% for claude-3.5-sonnet, and 78% for command-r-plus. This stereotype bias towards Baby Boomers may indicate that a small volume of training data is available where older generations are portrayed as liberal or neutrally affiliated.
Analyzing the outputs by gender, we see that gpt-4o-mini, claude-3.5-sonnet, and llama-3.1-70b all overwhelmingly select liberal (98%-100% of the time) for both males and females, suggesting no bias towards the selection of political affiliation based on gender. However, command-r-plus is much more likely to select conservative for males (70%) than for females (18%), suggesting that there is a gender bias that associates males with conservative politics.
Breaking down the outputs by ethnicity and race, we see that for all ethnicities and races, gpt-4o-mini, claude-3.5-sonnet, and llama-3.1-70b again all select liberal an overwhelming majority of the time (92%-100%) and select neutral 0% of the time for all ethnicities and races. This suggests that these models treat each ethnicity and race equally, and there is no bias. However, command-r-plus exhibits signs of racial and ethnic bias by disproportionately associating White individuals with a conservative political affiliation. It assigns "conservative" to White individuals 60% of the time, while significantly lower rates are observed for other groups: Neutral (20%), Black (0%), Hispanic (16%), and Asian (2%). Notably, all other racial and ethnic groups are predominantly categorized as liberal, suggesting that the bias is specific to White individuals. Additionally, when given a neutral ethnicity or race, command-r-plus selects "conservative" 20% of the time, which is higher than most other groups. This could be influenced by the model’s bias toward White individuals, as they make up the majority of the U.S. population. If the model interprets "neutral" as randomly selecting an ethnicity, it may be more likely to assign White, though this remains speculative.
Religion
Implicit
(a) Aggregate religion results across all models for females only.
(b) Aggregate religion results across all models for Baby Boomers only.
(c) Aggregate religion results across all models for White and Black individuals.
Figure 3: The religious affiliation distributions for texts generated using implicit inputs.
Tables A9, A10, A11 and A12 report the results of religion outputs when asked with implicit prompts. It is demonstrated that all four models we investigated tend to generate "unaffiliated" and "Christian" as the response for a person’s religion. We observe a substantial percentage of "Christian" in several demographic groups, particularly among females and Baby Boomers, where the results compared to real-world statistics are all highly significant. For example, gpt-4o-mini generates 86.4% of "Christians" for female prompts. On the contrary, non-Christian religions, such as Hinduism, Judaism, and Islam, consistently exhibit low percentages or zeros in most cases. These patterns reflect biases in LLM training that favor Christianity while underrepresenting diversity in religious and cultural affiliations. Meanwhile, in all models, the texts for White individuals are overwhelmingly Jewish (with a percentage ranging from 72% to 82%), and prompts with Black names result in the highest Muslim proportion (with a percentage ranging from 40% to 50%).
Diving into the results, claude-3.5-sonnet indicates that Asian people have a higher proportion of Buddhists (28%). Neutral and Hispanic people are predominantly unaffiliated. When examining age trends, gpt-4o-mini suggests that older generations, particularly Baby Boomers, are more Christian, while younger generations, like Generation Z and Generation Alpha, lean towards being agnostic or unaffiliated. For llama-3.1-70b, Asian and Black people show a higher Muslim proportion (30% and 40%, respectively).
Explicit
We evaluate the religion distributions generated by the four investigated LLMs according to different input demographics. It is demonstrated that when asked to provide the religion of a person, each LLM tends to output unaffiliated, Buddhist, and Christian, with a small proportion of responses being Hindu, Jewish, and Muslim. The bias in religion distributions—favoring unaffiliated, Buddhist, and Christian groups—can be attributed to the larger volume of human-written information online, which is subsequently used to train LLMs. Specifically, for demographic groups separated by gender, we first observe that female texts are more likely to be Buddhist, Christian, or unaffiliated. For example, for claude-3.5-sonnet, the percentage of female Buddhists is 10%. Next, for race and ethnicity, all models favor Buddhism among Asian people. For example, for gpt-4o-mini, the Asian texts are overwhelmingly Buddhist (98%) compared to other ethnicities. Similarly, for texts generated by llama-3.1-70b, most of the Asian samples are Buddhist (94%), whereas the White, Black, and Hispanic samples are almost all Christian (with 84% White Christian, 100% Black Christian, and 100% Hispanic Christian). Next, we analyze the texts among age demographics, for claude-3.5-sonnet and command-r-plus, Baby Boomers are completely Christian, and the remaining age groups are completely unaffiliated. For command-r-plus, Millennials, Generation Z, and Generation Alpha have a higher proportion of unaffiliated texts.
Sexual Orientation
Implicit
(a) Aggregate results across all models for sexual orientation with implicit gender input.
(b) Aggregate results across all models for sexual orientation with implicit ethnicity input.
(c) Aggregate results across all models for sexual orientation with implicit age input.
Figure 4: The sexual orientation distributions for texts generated using implicit inputs.
Examining the sexual orientation outputs of each model when given implicit prompts, all four models overwhelmingly generate sexual minorities (homosexual, bisexual, etc.) as the response for an individual’s sexual orientation. The percentages of heterosexual responses among the demographic groups are consistently low or even zero. To illustrate, claude-3.5-sonnet outputs 46% homosexual and 54% bisexual responses for Hispanic prompts, with 0% heterosexual responses. Gpt-4o-mini also generates 20% homosexual and 80% bisexual responses for Generation Z individuals. These results suggest that LLMs tend to significantly overrepresent minority sexual orientations compared to real-world statistics, where the majority of individuals identify as heterosexual [19].
Looking more closely at the results, claude-3.5-sonnet, llama-3.1-70b, and command-r-plus exhibit a higher proportion of heterosexual responses for White (16%, 10%, and 20%) and Black (32%, 20%, and 12%) people compared to other racial and ethnic groups. Claude-3.5-sonnet, gpt-4o-mini, and llama-3.1-70b also show that female individuals have higher bisexual percentages (68.8%, 89.4%, 80.8%) compared to males, who have a higher proportion of homosexual responses (88.8%, 60.0%, 79.6%). Analyzing the results by age group, claude-3.5-sonnet, gpt-4o-mini, and command-r-plus all demonstrate the highest proportion of heterosexual responses for Baby Boomers (46%, 18%, and 14%). On the other hand, for llama-3.1-70b, the only age group that shows a heterosexual percentage greater than zero is Generation Alpha (10%).
Explicit
(a) Aggregate results across all models for sexual orientation with explicit gender input.
(b) Aggregate results across all models for sexual orientation with explicit ethnicity and race input.
(c) Aggregate results across all models for sexual orientation with explicit age input.
Figure 5: The sexual orientation distributions for texts generated using explicit inputs.
When given explicit prompts, the LLMs also tend to overrepresent minority sexual orientations (homosexual, bisexual, etc.). However, an exception exists when the models are asked about the sexual orientation of Baby Boomer individuals. The percentage of heterosexual responses for Baby Boomers is 100% for claude-3.5-sonnet, llama-3.1-70b, and command-r-plus, and 98% for gpt-4o-mini. This stereotype bias towards Baby Boomers indicates how a small volume of training data is available where older generations identify themselves as having minority sexual orientations.
Analyzing the outputs by gender, claude-3.5-sonnet, gpt-4o-mini, and llama-3.1-70b are all more likely to generate bisexual responses for females (100%, 100%, 94%) and homosexual responses for males (96%, 60%, 96%). Notably, every one of the prompts asking claude-3.5-sonnet to generate details for White and Black male individuals, as well as a majority of the prompts about Asian female individuals, resulted in refusals. However, claude-3.5-sonnet provided a higher percentage of homosexual and bisexual responses for Neutral prompts (46% and 54%), while outputting a greater percentage of heterosexual responses for Hispanic (76%) individuals. For command-r-plus, the White texts are far more likely to be heterosexual (90%) compared to the other ethnicities (40% Neutral heterosexual, 28% Black heterosexual, 50% Hispanic heterosexual, and 32% Asian heterosexual). Similarly, llama-3.1-70b provides a higher proportion of heterosexual responses when asked to generate texts for White people (36%), whereas the texts for the remaining ethnic groups are more likely to be homosexual and bisexual.
Breaking down the results by age group, for each model, the responses are more likely to be heterosexual for Baby Boomers (100% for claude-3.5-sonnet, llama-3.1-70b, and command-r-plus, and 98% for gpt-4o-mini) and Generation X (74% for claude-3.5-sonnet, 48% for gpt-4o-mini, 68% for llama-3.1-70b, and 94% for command-r-plus) individuals. Llama-3.1-70b outputs a higher percentage of heterosexual responses for Generation Alpha (80%) as well. On the other hand, gpt-4o-mini provides a much higher percentage of bisexual responses for Millennials (98%) and Generation Z (98%).
Socioeconomic Status
Implicit
claude-3.5-sonnet
Upper-class Middle-class Lower-class Refusal
Gender Male (n=500) $5.60^{***}$ $93.80^{***}$ $60^{***} 0.0$00.00
Female (n=500) $7.80^{***}$ $91.80^{***}$ $0.00^{***}$ $0.40$
Ethnicity/Race Neutral (n=50) $8.00$ $88.00^{***}$ $0.00^{***}$ $4.00$
White (n=50) $38.00^{**}$ $52.00$ $4.00^{***}$ $6.00$
Black (n=50) $14.00$ $70.00^{***}$ $12.00^{***}$ $4.00$
Hispanic (n=50) $28.00^{***}$ $72.00^{**}$ $^{***} 0.000.00$
Asian (n=50) $42.00^{**}$ $48.00$ $0.00^{***}$ $10.00$
Age Baby Boomer (n=50) $6.00$ $94.00^{***}$ $^{***} 0.000.00$
Generation X (n=50) $2.00^{***}$ $98.00^{***}$ $^{***} 0.000.00$
Millennial (n=50) $2.00^{**}$ $98.00^{***}$ $^{***} 0.000.00$
Generation Z (n=50) $8.00$ $92.00^{***}$ $^{***} 0.000.00$
Generation Alpha (n=50) $22.00$ $78.00^{***}$ $^{***} 0.000.00$
Table 1: Socioeconomic status analysis of implicit bias for claude-3.5-sonnet.
command-r-plus
Upper-class Middle-class Lower-class Refusal
Gender Male (n=500) $42.20^{***}$ $57.80^{*}$ $^{***} 0.000.00$
Female (n=500) $31.80^{***}$ $68.20^{***}$ $^{***} 0.000.00$
Ethnicity/Race Neutral (n=50) $16.00$ $84.00^{***}$ $^{***} 0.000.00$
White (n=50) $20.00$ $80.00^{***}$ $^{***} 0.000.00$
Black (n=50) $22.00^{**}$ $78.00^{***}$ $^{***} 0.000.00$
Hispanic (n=50) $16.00$ $84.00^{***}$ $^{***} 0.000.00$
Asian (n=50) $42.00^{*}$ $58.00$ $^{***} 0.000.00$
Age Baby Boomer (n=50) $24.00$ $76.00^{***}$ $^{***} 0.000.00$
Generation X (n=50) $22.00$ $78.00^{***}$ $^{***} 0.000.00$
Millennial (n=50) $44.00^{***}$ $56.00$ $^{***} 0.000.00$
Generation Z (n=50) $30.00^{**}$ $70.00^{*}$ $^{***} 0.000.00$
Generation Alpha (n=50) $36.00^{***}$ $62.00$ $2.00^{***}$ $0.00$
Table 2: Socioeconomic status analysis of implicit bias for command-r-plus.
We examine the socioeconomic status outputs of each model when given implicit prompts. Notably, many models output a higher percentage of lower-class responses for Black people compared to other races. For example, when asked to provide the socioeconomic status of a Black person, 12% of claude-3.5-sonnet’s responses are lower-class, compared to 4% for White people and 0% for all other racial groups. When asked the same question with an Asian name in the prompt, claude-3.5-sonnet and command-r-plus both output a higher percentage of upper-class texts compared to other ethnic and racial groups. These differences may be attributed to how the LLMs’ training data are more likely to portray Black people as poor and Asian people as wealthy.
Diving deeper into the results, command-r-plus is more likely to provide upper-class responses for males (42.2%) compared to females (31.8%). The models also exhibit differences in the socioeconomic status distribution of their responses based on age group. For example, command-r-plus provides a lower percentage of upper-class responses for Baby Boomers (24%) and Generation X (22%), while gpt-4o-mini provides a higher percentage of middle-class responses for Millennials (78%) and Generation Z (82%). Moreover, claude-3.5-sonnet generates a higher proportion of upper-class texts for Generation Alpha (22%).
Explicit
claude-3.5-sonnet
Upper-class Middle-class Lower-class Refusal
Gender Male (n=50) $18.00$ $82.00^{***}$ $^{***} 0.000.00$
Female (n=50) $6.00$ $94.00^{***}$ $^{***} 0.000.00$
Ethnicity/Race Neutral (n=50) $8.00$ $92.00^{***}$ $^{***} 0.000.00$
White (n=50) $^{***} 0.000.00^{***}$ $0.00^{***}$ $100.00$
Black (n=50) $42.00^{***}$ $8.00^{**}$ $0.00^{***}$ $50.00$
Hispanic (n=50) $0.00^{*}$ $96.00^{***}$ $0.00^{***}$ $4.00$
Asian (n=50) $8.00$ $44.00^{***}$ $0.00^{**}$ $48.00$
Age Baby Boomer (n=50) $38.00^{***}$ $62.00$ $^{***} 0.000.00$
Generation X (n=50) $12.00^{*}$ $88.00^{***}$ $^{***} 0.000.00$
Millennial (n=50) $0.00^{***}$ $100.00^{***}$ $^{***} 0.000.00$
Generation Z (n=50) $0.00^{**}$ $100.00^{***}$ $^{***} 0.000.00$
Generation Alpha (n=50) $86.00^{***}$ $14.00^{***}$ $^{***} 0.000.00$
Table 3: Socioeconomic status analysis of explicit bias for claude-3.5-sonnet.
gpt-4o-mini
Upper-class Middle-class Lower-class Refusal
Gender Male (n=50) $2.00^{**}$ $84.00^{***}$ $14.00^{*}$ $0.00$
Female (n=50) $0.00^{***}$ $96.00^{***}$ $4.00^{***}$ $0.00$
Ethnicity/Race Neutral (n=50) $0.00^{***}$ $94.00^{***}$ $6.00^{***}$ $0.00$
White (n=50) $6.00^{**}$ $94.00^{***}$ $^{***} 0.000.00$
Black (n=50) $0.00^{*}$ $86.00^{***}$ $14.00^{***}$ $0.00$
Hispanic (n=50) $0.00^{*}$ $62.00$ $38.00$ $0.00$
Asian (n=50) $12.00^{*}$ $88.00^{***}$ $^{***} 0.000.00$
Age Baby Boomer (n=50) $0.00^{***}$ $100.00^{***}$ $^{***} 0.000.00$
Generation X (n=50) $0.00^{***}$ $100.00^{***}$ $^{***} 0.000.00$
Millennial (n=50) $0.00^{***}$ $100.00^{***}$ $^{***} 0.000.00$
Generation Z (n=50) $0.00^{**}$ $98.00^{***}$ $2.00^{***}$ $0.00$
Generation Alpha (n=50) $4.00$ $94.00^{***}$ $2.00^{***}$ $0.00$
Table 4: Socioeconomic status analysis of explicit bias for gpt-4o-mini.
When provided with explicit prompts, each of the models exhibits variations in the socioeconomic status distributions of the generated texts. For example, claude-3.5-sonnet generates a higher percentage of upper-class responses for males (18%) compared to females (6%). On the other hand, the texts generated by gpt-4o-mini are more likely to portray males as lower-class (14%) than females (4%).
Also, the differences exist among racial groups. For example, claude-3.5-sonnet outputs a higher proportion of upper-class responses for Black people (42%), followed by Neutral and Asian people (both 8%). However, command-r-plus outputs a far higher proportion of upper-class responses for White people (90%) and the Neutral group (82%), while gpt-4o-mini outputs a higher proportion of lower-class responses for Hispanic (38%) and Black (14%) individuals. Notably, claude-3.5-sonnet refused to generate texts for any prompts involving White individuals and Black males, as well as the majority of prompts concerning Asian females.
Nevertheless, many of the models still share similar types of stereotype bias. gpt-4o-mini and llama-3.1-70b are more likely to provide upper-class responses for Asian people. Generation Alpha also has the largest percentage of upper-class texts among age groups for claude-3.5-sonnet (86%), gpt-4o-mini (4%), and llama-3.1-70b (22%). Claude-3.5-sonnet and command-r-plus also output a higher percentage of upper-class responses for Baby Boomers (38% and 30%, respectively), compared to the other age groups.
Occupation
Implicit
llama-3.1-70b
Most Popular Occupations
Gender Male teacher (54.6%), software engineer (19.8%), graphic designer (16.8%), writer (4.0%), marketing specialist (1.4%)
Female teacher (52.2%), graphic designer (18.4%), marketing specialist (14.0%), social worker (5.0%), writer (2.6%)
Ethnicity/Race Neutral graphic designer (38.0%), teacher (34.0%), software engineer (20.0%), environmental scientist (2.0%), writer (2.0%)
White rabbi (24.0%), social worker (24.0%), writer (20.0%), software engineer (14.0%), teacher (6.0%)
Black writer (22.0%), graphic designer (22.0%), midwife (10.0%), software engineer (8.0%), community organizer (8.0%)
Hispanic teacher (50.0%), graphic designer (42.0%), writer (4.0%), urban planner (2.0%), event planner (2.0%)
Asian graphic designer (26.0%), monk (20.0%), journalist (14.0%), software engineer (14.0%), writer (8.0%)
Age Baby Boomer teacher (70.0%), software engineer (6.0%), writer (6.0%), social worker (6.0%), graphic designer (4.0%)
Generation X graphic designer (54.0%), teacher (22.0%), marketing specialist (10.0%), social worker (6.0%), writer (4.0%)
Millennial graphic designer (54.0%), marketing specialist (16.0%), software engineer (14.0%), teacher (8.0%), writer (4.0%)
Generation Z graphic designer (76.0%), teacher (10.0%), software engineer (8.0%), marketing specialist (2.0%), writer (2.0%)
Generation Alpha graphic designer (58.0%), writer (20.0%), lawyer/politician (6.0%), event planner (6.0%), lawyer (4.0%)
Table 5: Table analyzing implicit occupation bias statistics for llama-3.1-70b.
We first examine the occupation outputs of the models when given implicit prompts. A notable trend exhibited by all four models is that the most popular occupation provided for Baby Boomers is "teacher", which comprises 40.0% of the Baby Boomer responses for claude-3.5-sonnet, 44.4% for gpt-4o-mini, 70.0% for llama-3.1-70b, and 30.0% for command-r-plus. The occupations "teacher" and "graphic designer" are also present among the top five occupations for almost every demographic group in the texts generated by gpt-4o-mini, llama-3.1-70b, and command-r-plus. On the other hand, "engineer" is the most popular occupation for almost every demographic group for claude-3.5-sonnet. In addition to these trends, each of the models displays notable variations in the distribution of occupation outputs between gender, ethnicity, and age groups.
When analyzing the results by gender, the overwhelming majority of occupation responses for males given by claude-3.5-sonnet is "engineer" (95.2%). However, the occupation responses for female prompts are more evenly split, with "teacher" (24.4%), "engineer" (22.2%), and "executive" (21.4%) being the top three. Gpt-4o-mini is more likely to portray females as "social worker[s]" (33.0%), while llama-3.1-70b is more likely to output "software engineer" for males (19.8%) and "marketing specialist" for females (14.0%). Furthermore, command-r-plus is more likely to provide "financial analyst" (31.6%) and "teacher" (18.8%) as responses for males, whereas "social worker" (26.8%) and "artist" (14.2%) are the most popular occupations for females. These results suggest that each of the evaluated LLMs may hold unique occupation stereotypes towards users based on the gender of their name.
Looking at the LLM-generated texts for each ethnic group, both llama-3.1-70b (24.0%) and command-r-plus (16.0%) are more likely to provide "rabbi" as a response for prompts that included White names. Additionally, llama-3.1-70b outputs "midwife" for 10% of the Black prompts and "monk" for 20% of the Asian prompts. The other models also tend to provide a higher proportion of particular occupations for specific ethnic groups. For claude-3.5-sonnet, the second most popular occupation for Asian individuals is "executive" (20.0%), and for command-r-plus, the second and third most popular occupations for Hispanic individuals are "financial analyst" and "investment banker."
Explicit
gpt-4o-mini
Most Popular Occupations
Gender Male teacher (54.0%), software engineer (24.0%), graphic designer (18.0%), social worker (4.0%)
Female environmental scientist (44.0%), graphic designer (24.0%), social worker (16.0%), teacher (4.0%), community organizer (4.0%)
Ethnicity/Race Neutral teacher (42.0%), environmental scientist (24.0%), social worker (12.0%), graphic designer (6.0%), software developer (6.0%)
White software engineer (30.0%), marketing manager (28.0%), teacher (26.0%), project manager (8.0%), graphic designer (8.0%)
Black community organizer (64.0%), social worker (22.0%), teacher (10.0%), community outreach coordinator (4.0%)
Hispanic community organizer (48.0%), community health worker (20.0%), construction foreman (10.0%), social worker (8.0%), mechanic (4.0%)
Asian software engineer (92.0%), graphic designer (4.0%), social worker (2.0%), software developer (2.0%)
Age Baby Boomer teacher (90.0%), retired school principal (4.0%), retired teacher (4.0%), retired schoolteacher (2.0%)
Generation X marketing manager (38.0%), project manager (34.0%), graphic designer (16.0%), software developer (6.0%), software engineer (2.0%)
Millennial digital marketing specialist (72.0%), software developer (6.0%), graphic designer (6.0%), digital marketing manager (6.0%), marketing manager (6.0%)
Generation Z social media manager (58.0%), digital marketing specialist (20.0%), graphic designer (12.0%), freelance graphic designer (6.0%), sustainability consultant (4.0%)
Generation Alpha student (38.0%), software developer (16.0%), digital content creator (12.0%), digital marketing specialist (8.0%), content creator (4.0%)
Table 6: Table analyzing explicit occupation bias statistics for gpt-4o-mini.
When provided with explicit prompts, the LLMs also output varying occupation distributions based on the gender, age, or ethnic group mentioned in the prompt. The most popular occupation for Baby Boomers, "teacher," is the same across all four models (46% for claude-3.5-sonnet, 90% for gpt-4o-mini, 38% for llama-3.1-70b, and 26% for command-r-plus). Notably, for many of the models, the occupations common among the Generation Alpha texts are unique or have higher proportions compared to the other age groups. For example, popular Generation Alpha occupations include "student" (38.0%) and "digital content creator" (12.0%) for gpt-4o-mini; "student" (68%), "robotics engineer" (10%), and "environmental scientist" (4%) for llama-3.1-70b; and "entreprenuer" (20%), "social media influencer" (16%), and influencer (12.0%) for command-r-plus.
Each model provides different popular occupations for male and female prompts. To illustrate, for gpt-4o-mini, "teacher" (54.0%) and software engineer (24.0%) are the most popular occupations for males, while "environmental scientist" (44.0%) and "graphic designer" (24.0%) are the most popular occupations for females. On the other hand, command-r-plus’s most popular male occupations are "lawyer" (30.0%), "finance" (16.0%), and "financial analyst" (12.0%), while the most popular female occupations are "ceo" (18.0%), "prima ballerina (10.0%), and "executive" (10.0%).
When breaking down the results by ethnic group, both gpt-4o-mini and llama-3.1-70b are more likely to output "software engineer" as an occupation for White (30.0%, 48.0%) and Asian (92.0%, 42.0%) individuals. Llama-3.1-70b is more likely to portray Asian individuals as working in medical fields, with four of the top five occupations for Asians being dentist (46.0%), cardiologist (8.0%), pediatrician (2.0%), and dermatologist (2.0%). On the other hand, the most popular occupations for Black and Hispanic individuals are "community organizer" for gpt-4o-mini (64.0% for Black, 48.0% for Hispanic) and "teacher" for llama-3.1-70b (74.0% for both Black and Hispanic).
Polarity
Implicit
gpt-4o-mini
Median Standard Deviation Refusal
Gender Male 0.11 0.04 0.00
Female 0.14 0.04 0.00
Ethnicity/Race Neutral 0.14 0.05 0.00
White 0.13 0.05 0.00
Black 0.15 0.06 0.00
Hispanic 0.12 0.04 0.00
Asian 0.12 0.05 0.00
Age Baby Boomer 0.13 0.06 0.00
Generation X 0.13 0.05 0.00
Millennial 0.14 0.04 0.00
Generation Z 0.14 0.04 0.00
Generation Alpha 0.14 0.04 0.00
Table 7: Table analyzing implicit polarity bias statistics for gpt-4o-mini.
We perform sentiment score analysis for each generation. For each LLM-generated text, the polarity (i.e., sentiment score) is a value between -1, being the most negative, and 1, being the most positive. Overall, there exist few notable differences in sentiment scores between demographic groups. To illustrate, the female texts generated by gpt-4o-mini and command-r-plus have a slightly higher polarity value (0.14 and 0.15) compared to the male texts (0.11 and 0.12). For gpt-4o-mini and llama-3.1-70b, the Asian texts from both models also have the lowest median polarities (0.12 and 0.13) among all ethnic groups. This difference is especially notable for llama-3.1-70b, where the median polarity of the texts for all other ethnic groups is at least 0.17. Interestingly, some of the models exhibit contradictory sentiments towards the same demographic group. Claude-3.5-sonnet outputs the highest median polarity of 0.14 for Baby Boomer texts compared to other age groups. On the other hand, the Baby Boomer texts generated by command-r-plus have the lowest median polarity of 0.11, with all other age groups having a median polarity of at least 0.14.
Explicit
gpt-4o-mini
Median Standard Deviation Refusal
Gender Male 0.09 0.04 0.00
Female 0.13 0.05 0.00
Ethnicity/Race Neutral 0.11 0.05 0.00
White 0.08 0.05 0.00
Black 0.14 0.06 0.00
Hispanic 0.11 0.04 0.00
Asian 0.11 0.05 0.00
Age Baby Boomer 0.15 0.05 0.00
Generation X 0.08 0.06 0.00
Millennial 0.13 0.05 0.00
Generation Z 0.13 0.05 0.00
Generation Alpha 0.14 0.06 0.00
Table 8: Table analyzing explicit polarity bias statistics for gpt-4o-mini.
Looking at the sentiment scores among gender groups, the female texts generated by command-r-plus and gpt-4o-mini have higher median polarities of 0.22 and 0.13 compared to the male texts, whose median polarities are 0.18 and 0.09, respectively.
When analyzing the median polarity of the texts generated using explicit prompts, there exist multiple trends across gender and racial groups. The Black texts generated by claude-3.5-sonnet, gpt-4o-mini, and Llama-31.-70B have the highest median polarities (0.21, 0.14, and 0.18) compared to all other ethnic groups. For command-r-plus, the Black texts have the second highest median polarity of 0.18, while the Hispanic texts have the highest median polarity of 0.19. The Hispanic texts generated by llama-3.1-70b also have a higher median polarity of 0.17 relative to the other ethnic groups. On the other hand, the White texts generated by gpt-4o-mini and llama-3.1-70b have the lowest median polarities of 0.08 and 0.10, while claude-3.5-sonnet and command-r-plus have the lowest median polarities for Asian texts (0.10 for both).
Discussion
Large language models (LLMs) have the potential to increase human productivity in a plethora of domains, including medicine, education, law, and finance[1]. Considering the broad range of applications, it is essential to understand their potential risks and limitations. To investigate biases of LLMs, we provide implicit and explicit prompts for LLMs to generate profiles of individuals, depicting their political affiliation, religion, sexual orientation, socioeconomic status, and occupation. We then calculate the distribution of each demographic variable with respect to gender, race, or age. Our analyses show that the political affiliation, religion, sexual orientation, and socioeconomic status outputs of the LLM-generated texts exhibit significant stereotype and deviation biases for multiple groups. Following the evaluation procedure in the Model Evaluation Section in the Method Section, in summary, we report the stereotype and deviation biases for implicit and explicit prompts in Tables 910, respectively.
Politics. Regardless of prompt type, all four examined LLMs overrepresent individuals with liberal political affiliations. Despite frequently depicting people as liberal (sometimes 100% of the time), the LLMs rarely assign neutral affiliations (less than 10% in almost every case). Statistical tests confirm that these divergences from real-world political affiliation distributions are highly significant. As shown in Tables 910, the political deviation bias score is close to 1 for all models. For users, these findings highlight the importance of critical engagement with model outputs, especially when using LLMs for sociopolitical analysis or representation. In sociopolitical analysis, organizations should explicitly communicate the limitations of model outputs. This helps ensure that decisions are not made unknowingly based on biased decisions.
Religion. When analyzing the religious distributions of the LLM-generated texts, all four LLMs are most likely to portray individuals as being "Christian" or "unaffiliated". In particular, over 50% of Baby Boomers are labeled as Christian across all models and prompt types. These outputs likely reflect biases in the training data, which heavily feature English or U.S.-centric content about individuals from that age group. Additionally, the proportions of Hindu, Jewish, and Muslim texts are overwhelmingly small. Although the population of people adhering to these religions is not negligible in reality, LLMs show a consistent preference for the Christian and unaffiliated groups and sparsely select Hindu, Jewish, and Muslim texts. While the deviation bias score is less than 0.500 for all models and prompt types (except for llama-3.1-70b when given implicit prompts), the stereotype bias score for religion in these tables is very high, which is likely explained by the fact that religion is correlated with ethnicity (see Tables 910).
For LLM designers, the consistent overrepresentation of Christian and unaffiliated religious identities highlights the need to critically examine and diversify training data to reflect global religious and cultural distributions. For users, it is essential to approach model output with awareness that such responses may reflect underlying data biases rather than objective truths, particularly when generating content about identity-related attributes such as religion.
Sexual Orientation. Additionally, across both implicit and explicit prompts, all four models overrepresent sexual minorities compared to real-world statistics. For example, claude-3.5-sonnet, gpt-4o-mini, and llama-3.1-70b are more likely to portray females as bisexual and males as homosexual. Overall, heterosexual identities are underrepresented in LLM outputs for most age groups, revealing a consistent sexual orientation bias. As demonstrated in Tables 910, the deviation bias score for every model when given implicit prompts is 1, indicating that every output attribute is significantly different for sexual orientation. Additionally, the stereotype bias is above 1 for every model (except for command-r-plus).
The consistent overrepresentation of LGBTQ identities, especially among younger generations, highlights the need to inspect training data and reevaluate model alignment strategies to ensure they reflect real-world distributions more accurately. Additionally, providing greater transparency around how models are aligned would help users better understand the potential biases present in their outputs. While inclusivity is valuable, inflated portrayals unintentionally distort social realities, leading to skewed perceptions or reinforcement of stereotypes.
Socioeconomic Status. The four LLMs show notable variations in the socioeconomic status distributions of their generated texts as well, as depicted in Tables A25-A32. Regardless of prompt type, model, or demographic group, the majority of the texts’ socioeconomic status proportions are statistically significant when compared with real-world reference proportions. The deviation bias for both implicit and explicit prompts is similar, and claude-3.5-sonnet and gpt-4o-mini display high levels of stereotype bias compared to the other models. Moreover, some LLMs are more likely to provide upper-class responses for Asian individuals for both implicit (claude-3.5-sonnet and command-r-plus) and explicit (gpt-4o-mini and llama-3.1-70b) prompts.
Such a trend in the LLM-generated texts may conflict with the goal of providing equitable services to all users, no matter their racial identity. LLM developers should carefully weigh the tradeoffs between ensuring that LLM responses accurately portray reality and preventing models from perpetuating harmful stereotypes. This balance needs to be considered during both the LLM training and fine-tuning processes, when developers must evaluate what training data should be incorporated and what outcomes reinforcement learning from human feedback (RLHF) should prioritize, respectively.
Occupation. Looking at the occupation outputs of the LLMs, each model outputs a variety of occupations for each demographic group, with notable differences in the distribution of occupations among gender, racial, and generational lines. These differences exist among the texts generated using both implicit and explicit prompts. While the specific details are reported in the results section and Tables A33-A40, an overarching trend exists where all four models, regardless of prompt type, provide "teacher" as the most popular occupation for Baby Boomers.
Polarity. Considering the sentiment scores of LLM-generated texts, the texts generated using explicit prompts display a much larger variation of polarity scores between demographic groups. On the other hand, the texts generated using implicit prompts display few notable differences in sentiment scores between demographic groups, and the existing differences are also much smaller. This suggests that LLMs are more sensitive to bias when provided with prompts that contain explicit mentions of gender, ethnicity, or age groups. For example, when provided with explicit prompts, LLMs tend to generate texts with higher polarity scores for historically marginalized groups, such as Black, Hispanic, or female individuals. In contrast, the texts for White, Asian, or male individuals tend to have lower polarity scores.
These differences may be due to the possibility that the data used to train LLMs portrays historically marginalized groups in a more positive light. However, such an outcome may occur at the expense of White, Asian, and male individuals. LLM users should be wary of such discrepancies in responses when providing LLMs with prompts that explicitly mention a group or individual’s gender, race, or age. Furthermore, LLM developers should incorporate training data that represents a diverse range of sentiments for various demographic groups to ensure equitable responses. During the fine-tuning process, developers may also consider penalizing large variances in the sentiment scores of responses when given similar prompts by users from diverse ethnic backgrounds.
Refusal. Notably, claude-3.5-sonnet was the only model in our study that refused to answer any prompts. When provided with explicit prompts, claude-3.5-sonnet refused 100% of White, 50% of Black, 4% of Hispanic, and 48% of Asian prompts. For the explicit prompts involving Asian and Black individuals, all of the refusals were for Asian females and Black males, respectively.
Refusing to answer inappropriate prompts decreases the risk of an LLM generating harmful or dangerous content. While certain prompts, such as those asking a model to generate violent or offensive texts, are obviously inappropriate, the definition of what makes a prompt appropriate may be less clear in other cases. During the alignment and fine-tuning processes, LLM developers must carefully weigh the balance between effectively serving users’ needs and having guardrails in place to prevent a model from generating harmful content. Future research should investigate the effectiveness of various model alignments and fine-tuning techniques in achieving this balance.
Limitations and Future Work. Alongside the significance of our study’s findings, its limitations also present ideas for future research to more deeply explore the nature and extent of LLM biases. First, the input attributes used to prompt the models in our study are limited. In addition to using gender, race, and age as inputs, prompting models to depict users based on their religion, sexual orientation, or disability may lead to deeper understandings of how LLMs perceive various demographic groups. Furthermore, the ethnic groups used in our prompts represented White, Black, Hispanic, and Asian individuals. Investigating LLM outputs portraying more diverse ethnic groups (e.g., Middle Eastern or South Asian) may yield more insightful findings. Second, the real-world reference statistics used in our study to calculate deviation bias were for the United States population[20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. Subsequent studies may investigate the deviation bias of LLMs with respect to population statistics for other countries. Finally, our study asked LLMs to generate depictions of individuals belonging to certain gender, race, or age groups without additional information or context. Future research could explore how LLM outputs can differ for prompts with varying demographic inputs when provided with specific tasks based on real-world use cases, such as financial or career advice, medical diagnosis, or creative writing.
Implicit claude-3.5-sonnet gpt-4o llama-3.1-70b command-r-plus
Stereo. Dev. Stereo. Dev. Stereo. Dev. Stereo. Dev.
Politics 5.219 1.000 0.373 0.972 0.966 0.972 1.848 0.889
Religion 11.266 0.486 9.157 0.347 8.512 0.542 8.665 0.250
Sexual Orientation 7.958 1.000 1.782 1.000 1.976 1.000 0.137 1.000
Socioeconomic Status 1.394 0.806 0.652 0.583 2.439 0.944 0.095 0.778
Table 9: Implicit bias metrics across models. Each cell reports two values for a given demographic category and model: Stereotype Bias (Stereo.), computed as the mean of the maximum Kullback–Leibler divergences across gender, ethnicity, and age distributions  (1); and Deviation Bias (Dev.), calculated as the proportion of statistically significant p-values across pairwise comparisons within each category  (2). Higher values indicate greater bias.
Explicit claude-3.5-sonnet gpt-4o llama-3.1-70b command-r-plus
Stereo. Dev. Stereo. Dev. Stereo. Dev. Stereo. Dev.
Politics 20.063 0.944 2.100 0.972 11.440 0.972 14.379 0.778
Religion 20.854 0.431 21.078 0.292 11.798 0.319 1.484 0.264
Sexual Orientation 11.587 0.958 12.244 0.958 14.259 0.958 0.823 0.875
Socioeconomic Status 14.224 0.861 4.026 0.917 0.577 0.972 0.765 0.861
Table 10: Explicit bias metrics across models. Each cell reports two values for a given demographic category and model: Stereotype Bias (Stereo.), computed as the mean of the maximum Kullback–Leibler divergences across gender, ethnicity, and age distributions  (1); and Deviation Bias (Dev.), calculated as the proportion of statistically significant p-values across pairwise comparisons within each category  (2). Higher values indicate greater bias.
Methods
Data
Real-World Distributions
We chose three demographic groups: gender (male, female), ethnicity/race (White, Black, Hispanic, Asian, Neutral), and age (Baby Boomer, Generation X, Millennial, Generation Z, Generation Alpha). We obtained real-world data on the distributions of output attributes for various input attributes from a variety of reputable sources that provide demographic breakdowns for the U.S. population. These sources include government census data, national surveys, and peer-reviewed research that reports population percentages for each category in different demographic groups. Specifically, we relied on datasets that detail political leanings, religious affiliations, sexual orientation, and socioeconomic status from the Pew Research Center, Public Religion Research Institute, Gallup, Williams Institute, Springtide Research Institute, and Ipsos [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. For each of the output categories, we ensured that the data reflected national averages and demographic breakdowns between gender, ethnicity/race, and age. These real-world distributions served as a benchmark for evaluating the deviation bias in the LLM outputs.
Generated Distributions
Before generating the data, we first constructed a list of identifiers explicitly and implicitly linked to gender, ethnicity/race, and age to serve as input attributes for generating descriptions using LLMs. After selecting these input attributes, we designed a prompt template that would take explicit and implicit identifiers of these categories. This prompt template and example model output are shown in Figure 6(a) and Figure 6(b), respectively.
(a) Example prompt template.
(b) Example model output.
Figure 6: Prompt template and Model output example
For explicit identifiers, we filled in our prompt with the attributes within each category. For example, we ask the model, "Write a 200 word description of a Hispanic man." For implicit identifiers, we filled in our prompt with names that have a high correlation with these attributes to implicitly indicate such characteristics. For example, we ask the model, "Write a 200 word description of Juan." We selected such names accordingly. For gender, we selected the top 100 baby names for males and the top 100 baby names for females over the last 100 years in the United States based on data from the Social Security Administration [35]. For ethnicity/race, we chose the five most ethnically specific names for White, Black, Hispanic, and Asian males, females, and the general population [21]. To represent age, we found the five most popular male and female baby names that were unique to each generation. According to the database [36], the age groups are defined as: Baby Boomers (1946–1964), Generation X (1965–1980), Millennials (1981–1996), Generation Z (1997–2012), and Generation Alpha (2013–2023). For each age group, we selected the five most popular male and female baby names that were present among the top 200 most popular male and female names for the corresponding generation but not among the top 200 most popular names in any other generations [20, 36].
After gathering the implicit and explicit identifiers, to ensure a robust and reliable analysis, we performed five independent generations for each implicit prompt, 25 independent generations for each explicit prompt indicating ethnicity/race and age, and 50 independent generations for each explicit prompt indicating gender. These numbers were chosen to increase the sample sizes and capture variability across outputs, resulting in matching sample sizes for the same demographic groups between implicit and explicit prompts.
All models were set with a temperature of 0.7 and top_p of 0.9. The rest of the attributes were default, with the exception of Claude-3.5-Sonnet, where "max_tokens" was set to 1000. In each generation, we asked the model to provide an attributes section containing the attributes of the individual, which we name output attributes. These output attributes contain political affiliation (conservative, liberal, neutral), religion (Christian, Buddhist, Hindu, Jewish, Muslim, unaffiliated), sexual orientation (heterosexual, homosexual, bisexual), and socioeconomic status (lower-class, middle-class, upper-class). Some outputs from the models were not part of these categories, so we manually fit them to such categories (e.g., “lesbian” was changed to “homosexual”, "Catholic" was changed to "Christian"). For sexual orientation, the categories were then grouped into "heterosexual" and "LGBTQ" for ease of analysis with real-world population percentages. The generated attributes were verified to appear in the model’s descriptions through a combination of keyword searches and manual review.
With these text generations, we then counted the number of times the LLM generates each of the output categories for a given input attribute across the generations of outputs. For example, if we aim to analyze the political affiliation for the input attribute "Hispanic," we count how many times the model generates "conservative," "liberal," and "neutral" as political affiliations over all generations with the input category of Hispanic.
Investigated LLMs
We carefully selected four LLMs for our evaluation and chose these models based on parent company, consumer usage, and LLM’s ability to consistently generate text that followed our instructions and was easy to parse. Our chosen models include claude-3.5-sonnet from Anthropic [37], gpt-4o-mini from OpenAI [38], command-r-plus from Cohere [39], and llama-3.1-70b from Meta [40]. All models come from different companies, are the flagship models in those companies (as of August 2024)[37, 38, 39, 40], and followed our instructions carefully by including an "attributes" section within their output that accurately reflected the generated paragraph.
Model Evaluation
In evaluating the bias of each LLM, we carefully compare the LLM-generated distributions for each output category (e.g., religion) for each input attribute (e.g., male and female). We analyze the generated distributions in two main ways, which we call stereotype bias and deviation bias.
Stereotype Bias
To evaluate an LLM’s stereotype bias towards a specific demographic group, we compare each of the input attributes’ generated distributions with respect to a given output category. For example, to determine whether an LLM displays socioeconomic stereotype bias against a certain gender, we compare the socioeconomic status distributions of the LLM-generated texts for each gender. This analysis will tell us if a model treats a certain input attribute (e.g., male) differently from its complementary input attribute(s) (e.g., female).
Additionally, in Tables 9 and 10, we measure stereotype bias by first computing the maximum Kullback–Leibler divergence [41] (KL) between any pair of output categories within each attribute group: gender, ethnicity, and age. We then take the mean of these three maximum values to estimate the model’s overall stereotype bias. Higher KL divergence values indicate greater disparities between demographic groups, and thus suggest stronger potential for stereotype bias.
$\text{Stereotype Bias Score}=\text{mean}\left(\max\text{KL}_{\text{gender}},\max\text{KL}_{\text{ethnicity}},\max\text{KL}_{\text{age}}\right)$ (1)
Deviation Bias
For deviation bias, we compare the distribution of output attributes in the LLM-generated texts for a specific input demographic group with the real-world, ground truth distribution of attributes for that demographic group. This test is crucial because the real-world distribution of a certain demographic attribute (e.g., political affiliation) often depends on our chosen input attributes (e.g., age). The deviation bias metric complements the stereotype bias metric, providing a clearer view of the model’s true biases. To calculate the deviation bias and compare the demographic distributions of the LLM-generated texts with the real-world, ground truth demographic distributions, we perform a binomial test on the frequency of each output attribute. We assess whether the observed frequency for each attribute significantly deviates from what we expect based on the real-world frequency. This method enables us to account for subtle differences that might not be captured by the stereotype bias metric alone, ensuring that the evaluation reflects realistic variations rather than misinterpreting them as bias.
In Tables 9 and 10, we report a deviation bias metric, defined as the proportion of output attributes showing statistically significant differences from the real-world population (P < 0.05). Values closer to 1 indicate a higher number of significantly biased output attributes, reflecting a higher level of deviation bias for a given model.
[…]
Author Contributions
B.E., M.M., and W.D. conceived the study. W.D., B.E., M.M., and F.X. designed the study. W.D. and B.E. collected and preprocessed data. B.E. and W.D. analyzed data. W.D., B.E., M.M., and F.X. wrote the paper.
Data Availability
Data and codes used in this study are available at: https://github.com/daedaldan/llm-stereotype-deviation-biases.
Competing Interests
The authors declare no competing interests.
Funding
The authors declare that no funds were received.